# Geometric QCD II: The Confining Twistor String and Meson Spectrum

### 12 Mar 2026 v3

@@ 8 Asymptotically free planar QCD from the loop equation
The Elfin theory induces QCD in the limit of large fermion mass, serving as a short-distance cutoff. There are no singularities at finite mass, which eliminates the messy discussion about cusp and perimeter singularities of the Wilson loop. The Wilson loop is defined in the Elfin theory as a regularized solution of the loop equation, with the delta function smeared by a finite fermion mass. This definition of the loop equation makes mathematical sense, in our opinion: regularize this equation by smearing the delta function on the right side while preserving the gauge invariance by keeping all loops closed. The Elfin theory falls into the category of regularized solutions of the loop equation.
@@ 8.2 Gluon graphs by iterations of Bootstrap equation
The derivation was rather tedious, but it faithfully reproduced conventional planar graphs, including Faddeev-Popov ghost loops. The first order planar graph follows immediately from the (8.1), as shown on Fig. 11, by replacing $W$ by $1$ in the path integral, after which this path integral yields the gluon propagator $1/(x-y)^{2}$. The second order graphs are already nontrivial (see [25], Appendix B). This is where the Faddeev-Popov ghost loops emerge directly from iterations of the bootstrap loop equation; this is also where the beta function appears, leading to asymptotic freedom with its running coupling constant. Perturbative QCD is reproduced by the bootstrap equation to all orders, as it was further investigated in [23].
@@ 9 Effective action, its geometry and its singularities
Where is this non-locality hiding in the Elfin theory, and how will these logarithms emerge? In this section, we dig deeper into the formal solution we have found, searching for the perturbative singularities of planar QCD.
@@ 9.1 Local expansion of the heavy fermion determinant
The leading term is the Area Law, where the tension $\sigma\sim m^{2}$ provides the confining potential. The subleading term is the conformal anomaly (Liouville action). There is always another zero mode confining factor $G[C]=\exp\left(-\delta\sigma S[C]\right)$, renormalizing the string tension $\sigma$, as we found in the previous work [36]. Equivalently, the area term generated in the heavy-mass expansion of the determinant provides an additive renormalization of the vacuum-energy/string tension sector, already parametrized by the zero-mode dressing $\exp\left((\right)-\kappa S[C])$; only the sum defines the single physical string tension $\sigma_{phys}=2\sqrt{2}\kappa_{phys}$ in the continuum theory, while $M$ acts strictly as a regulator scale removed in the local limit.
@@ 9.3 New view at the cusp singularity
This mechanism is strictly analogous to the continuum limit in Lattice Gauge Theory, where the bare coupling $g_{0}\to 0$ as the lattice spacing $a\to 0$ to keep physical masses finite. In our induced QCD, the cutoff has a physical and mathematical origin in the regularization of the minimal surface, so the limit $m\to\infty$ (where $\lambda\to 1/(\beta_{0}\log m)$) has a precise mathematical meaning, replacing the formal subtraction of infinities with a finite limiting process.
@@ 10.1 Regularization, renormalizability, and the microscopic definition of Planar QCD
However, the lattice definition comes at a cost: it explicitly breaks the continuous symmetries of four-dimensional Euclidean space, from the rotation group $O(4)$ to subtle topological structures like the Hodge duality of the area tensor. For instance, exact instanton solutions or Hodge-dual minimal surfaces do not exist on a discrete grid; they only emerge in the continuum limit.
@@
History offers cautionary tales. The large-$N$ reduced Eguchi-Kawai model, for example, encountered unphysical fixed points. Similarly, the original Induced QCD proposal [20]—which attempted to induce gauge dynamics via heavy free fermions with vanishing color currents on a lattice—failed because the theory collapsed into the wrong phase rather than asymptotically free QCD. Our Elfin theory is a variation on this theme: heavy free fermions inducing QCD. However, the critical distinction is that our fermions live on a rigid minimal surface rather than a four-dimensional lattice. This geometric constraint keeps the theory non-trivial in the local limit, making it a viable regulator for Planar QCD.
A central question arises regarding the renormalization of the Wilson loop itself. Standard perturbation theory dictates that $W[C]$ requires multiplicative renormalization to remove divergences associated with cusps and self-intersections. Are we obliged to redefine the loop equation to make the coordinate-space Wilson loop finite in the local limit?
We argue that this is a misguided requirement that obscures the underlying dynamics. Physical renormalizability demands only the existence of a universal mass scale $\Lambda_{QCD}$ and the finiteness of observable scattering amplitudes. The Wilson loop $W[C]$ is an intermediate, off-shell quantity. Physical observables involve path integrals over quark trajectories, which are governed by a Brownian measure. A Brownian trajectory is nowhere differentiable; it consists of an infinite density of cusps. Attempting to renormalize the Wilson loop for a smooth contour $C$ is physically artificial because the dominant contributions to the path integral come from fractal paths where the concept of a "cusp angle" is ill-defined.
@@ 10.2 Momentum loop path integral
As we discussed in the previous section, the ordinary Wilson loop is not directly observable, and its notorious singularities are merely intermediate results that will eventually disappear (cancel or integrate out) in observables.
Unlike the standard Feynman-Schwinger worldline representation in coordinate space, which is inherently plagued by the UV cusp singularities of $W[C]$, we utilize the phase-space path integral representation introduced by the author in [30]. In this formulation, the exact planar quark amplitude factorizes strictly into a universal kinematic Dirac phase-space trace $\mathcal{K}[P]$ and the dynamical gluonic momentum-loop functional $W[P+Q]$.
@@
To appreciate the mathematical control within this phase-space formulation, it is instructive to examine its exact canonical normalization. The functional measure is strictly normalized by the standard symplectic volume element $\mathcal{D}P\mathcal{D}C^{\prime}/(2\pi)$ per continuous degree of freedom. Consequently, in the absence of the interacting Wilson loop (the free-particle limit, $W[C]=1$), integrating over the loop coordinates $C$ natively yields a functional Dirac delta function, strictly enforcing the conservation of local momentum:
@@
This demonstrates that the phase-space integral is rigorously well-defined from standard quantum mechanics, unique strictly modulo the infinite 1D diffeomorphism volume $\text{Vol}(\text{Diff}(S^{1}))$ associated with the simultaneous reparametrization of $P(\tau)$ and $C(\tau)$.Conventionally, this $\text{Diff}(S^{1})$ volume is factored out by imposing a proper-time gauge-fixing via an $\int(C^{\prime})^{2}d\tau$ kinetic term. This specific gauge choice inevitably generates the Feynman-Schwinger representation with its standard Brownian (Wiener) measure in loop space—a measure inherently dominated by everywhere-discontinuous fractal paths that artificially generate the notorious UV cusp singularities. In our framework, we recognize that the Brownian measure is just one possible choice of gauge fixing for the 1D diffeomorphisms. We will deliberately abandon it. Instead, as detailed in Section 11, we fix $\text{Diff}(S^{1})$ by imposing the Virasoro constraint $(f^{\prime})^{2}=0$ on the holomorphic twistors parametrizing the loop velocity. This specific gauge choice entirely avoids the fractal artifacts of the Brownian measure and perfectly locks the 1D phase-space kinematics of the quark loop to the 2D internal confining geometry of the minimal surface.
@@ 10.3 The Momentum loop equation
The notorious delta function on the right (which caused all the perturbative singularities in $W[C]$) disappeared in the Fourier integral (see original papers [29, 30, 31] for details).
@@ 10.4 Equivalence of loop diffusion operator to the third Magnus form
where $\Theta(\theta-\tau)$ denotes the periodic theta function on the circle, the trilinear product is transformed into a general triple integral of the momentum derivative $P^{\prime}$. The rigorous mathematical foundation for expanding loop functionals in such iterated integrals is provided by Chen’s theorem [1]. Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a strictly complete functional basis. This guarantees that our formal expansion in terms of Magnus forms is exhaustive.
Through direct tensor contraction, the kinematic tensor $T^{\alpha\beta\gamma}{}_{\nu}$ implements the exact Dynkin projector [9] onto the nested double commutator $\hat{D}_{\alpha},[\hat{D}_{\alpha},\hat{D}_{\nu}]]$. Here, $\hat{D}=\partial+A$ represents the covariant derivative operator within the gauge field representation of the Wilson loop. This mapping identifies the associative word structure of the integrated momentum field with its unique Lie algebra representation.
Consequently, by Ree’s Theorem [43] for Free Lie Algebras [44], this operator identically annihilates the symmetric shuffle ideal associated with the triple integral. As the point-split coordinates $\theta_{i}$ collapse to the origin, the periodic step-functions define a strictly cyclically ordered simplex. This unique parametric-invariant Lie projection ensures that the functional converges exactly to the third-order Magnus form $\Omega^{(3)}$. Direct evaluation of this triple integral, mediated by the $T$ projector, confirms this algebraic correspondence and yields the canonical normalization coefficient $1/6$.
@@ 10.5 Algebraic recurrence for the Momentum Loop Equation
The Momentum Loop Equation (MLE) exhibits a profound simplification when the momentum functional is represented as a Magnus-ordered exponential:
@@
where $\hat{X}_{\mu}$ represents the operator position of the string endpoint in spacetime. Expanding this trace in Magnus forms $\Omega^{n}$, we obtain universal tensor coefficients $\mathcal{W}_{\alpha_{1}\dots\alpha_{n}}^{(n)}=\mathrm{tr}\{\hat{X}_{\alpha_{1}}\dots\hat{X}_{\alpha_{n}}\}$. The functional derivative acts algebraically on this structure, bringing down exact commutator insertions:
@@ 10.6 Exact solution up to $\mathcal{O}(P^{\prime 7})$ and non-analyticity of the Taylor expansion
When substituted into the iterated integrals of the Magnus expansion, the $\delta$-function explicitly inserts the macroscopic gap vector $\Delta P_{\mu}\equiv\Omega^{(1)}_{\mu}$ precisely at the boundaries of the integration domain. Because $\Delta P$ corresponds to the first-order Magnus form, this topological closure modifies the $n$-th order form by subtracting the product $\Omega^{(1)}\Omega^{(n-1)}$. By the integration-by-parts identity of the Magnus expansion (the shuffle relations), this product translates exactly to the sum of all interleaved operator positions. The loop-closing transformation thus yields the modified Magnus form:
@@
Substituting this closed-loop basis into the MLE, we dynamically equate the left-hand side kinematic shuffle products against the right-hand side discrete commutator series. Starting from the perturbative vacuum $\mathcal{W}^{(0)}=1$, the algebraic system successfully yields finite, exact solutions for the lowest-order tensors. The fundamental string tension enters at $\mathcal{W}^{(2)}$, and the kinematic stress generated by the loop derivative exactly constrains the cyclic symmetric pairings of $\mathcal{W}^{(4)}$ and $\mathcal{W}^{(6)}$. (The explicit analytic formulas for these invariant tensors are provided in C for small $n$. Crucially, C also contains the exact finite-dimensional linear algebra diagnostics (Table 1) and the explicit Fredholm inconsistency proof demonstrating the absolute lack of solutions at $\mathcal{W}^{(8)}$. The general algorithmic implementation is provided in [34].)
@@ 10.7 The Reasons for the Failure of the Magnus Expansion for $W[P]$
The general reason for the breakdown of the expansion at this order is that the linear space of $\mathcal{W}^{(8)}$ invariant tensors—which is strictly constrained by the Shuffle Ideal and cyclic symmetry—possesses a significantly smaller matrix rank than the highly specific multilinear stress generated by the crossing Kronecker deltas of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$.
This is not a heuristic observation, but a rigorously quantified algebraic contradiction. As detailed in C (and summarized in Table 1), extracting the exact linear system at the 8th order yields 379 geometric constraints but only 212 available parameters (which exhaustively includes the full $\mathcal{W}^{(8)}$ basis, all residual lower-order free parameters, and all gauge/ideal freedoms). The resulting system explicitly violates the Fredholm alternative condition, yielding a strictly positive rank deficiency ($\Delta\text{rank}=1$) and a non-zero projection of the inhomogeneous stress onto the left-null space of the kinematic matrix.
@@
The physical meaning of this failing Taylor–Magnus expansion is of paramount importance. It demonstrates that the MLE is finite, free of UV divergences, and expandable in a functional Taylor series up to $\mathcal{O}(P^{\prime 6})$. However, the geometric contradiction at the 8th order rigorously proves that the exact solution $W[P]$ is not an analytic functional of the bounding loop $P(\theta)$. It cannot be indefinitely expanded in a continuous Taylor–Magnus series.
@@
governing a scalar functional $W[P]$. The number of independent tensor structures available for a scalar functional grows strictly slower than those required for a free-index vector equation (growing approximately four times slower in $d=4$). This is why, at some critical level of the Magnus expansion, there will inevitably be more vector equations than available scalar parameters.
This is a foundational feature of the loop equation: it descends directly from the vector Yang–Mills equations of motion, which dictate that a local vector loop equation (evaluated at a fixed point on the contour) must be satisfied. Resolving this overdetermined system demands a deep mathematical “conspiracy” among the internal degrees of freedom within $W[P]$. The generic Taylor–Magnus expansion simply lacks the internal geometric structure to orchestrate this conspiracy.
(Interestingly, this exact same vector-versus-scalar mismatch is fatal to all naive bosonic string theories of QCD: the loop equations are strictly vector equations, but the Nambu–Goto string Laplace operator is a scalar, meaning a scalar area-derivative constraint fundamentally cannot capture the full vector loop equation without an anomaly.)
However, our theory natively provides this required conspiracy! In the previous sections, we derived exact vector equations for the Elfin determinant on the Hodge-dual minimal surface, and these vector equations are satisfied across all vector components. This mathematical miracle is made possible by Hodge duality and the planar factorization of the Majorana determinant on this Hodge dual surface.
The profound physical reasons for this non-analyticity will become fully apparent in later sections, when we construct the continuum solution and reduce it to the exact twistor string path integral. In that formulation, the formal Taylor expansion in $P(\theta)$ corresponds to the perturbative expansion of the exact exponential phase factor:
@@
For such a functional Taylor series to converge, the boundary twistors $\lambda(z)$ and $\mu(z)$ would need to be strictly bounded. The algebraic contradiction at the 8th order rigorously proves this is not the case; the path integral is not convergent in a perturbative Taylor expansion.
As we will discuss in detail when deriving the twistor string path integral, the integration over the momentum loops is not dominated by smooth, small $P$ variations at all. When integrated alongside the Dirac target-space product $\mathrm{tr}\,\hat{\mathcal{T}}\exp\left(-\mathrm{i}\int d\theta\,\gamma_{\mu}P_{\mu}(\theta)+m_{0}\right)$, we encounter strictly ultra-local integrals. At every angle $\theta$, we integrate out the local value $P(\theta)$. Because there are no worldline kinetic derivatives $\partial_{\theta}P$ in the action, the path integral completely factorizes point-by-point. As a result, the dominant trajectories are neither continuous nor dominated by small $P$ variations. We therefore conclude that the MLE is completely finite and anomaly-free, but its true continuum solution is inherently non-perturbative and cannot be Taylor-expanded beyond the 6th order.

### 23 Mar 2026 v4

@@ 8 Asymptotically free planar QCD from the loop equation
The Elfin theory induces QCD in the limit of large fermion mass, serving as a short-distance cutoff. There are no singularities at finite mass, which resolves the ambiguities associated with cusp and perimeter singularities of the Wilson loop. The Wilson loop is defined in the Elfin theory as a regularized solution of the loop equation, with the delta function smeared by a finite fermion mass. This naturally regularizes the loop equation by smearing the delta function on the right side while preserving the gauge invariance by keeping all loops closed. The Elfin theory falls into the category of regularized solutions of the loop equation.
@@ 8.2 Gluon graphs by iterations of Bootstrap equation
Iterative evaluation reproduces conventional planar graphs, including Faddeev-Popov ghost loops. The first order planar graph follows immediately from the (8.1), as shown on Fig. 11, by replacing $W$ by $1$ in the path integral, after which this path integral yields the gluon propagator $1/(x-y)^{2}$. The second order graphs are already nontrivial (see [21], Appendix B). This is where the Faddeev-Popov ghost loops emerge directly from iterations of the bootstrap loop equation; this is also where the beta function appears, leading to asymptotic freedom with its running coupling constant. Perturbative QCD is reproduced by the bootstrap equation to all orders, as it was further investigated in [19].
@@ 9 Effective action, its geometry and its singularities
We now trace the origin of this non-locality and the emergence of logarithmic singularities within the Elfin theory.
@@ 9.1 Local expansion of the heavy fermion determinant
The leading term is the Area Law, where the tension $\sigma\sim m^{2}$ provides the confining potential. The subleading term is the conformal anomaly (Liouville action). There is always another zero mode confining factor $G[C]=\exp\left(-\delta\sigma S[C]\right)$, renormalizing the string tension $\sigma$, as we found in the previous work [32]. Equivalently, the area term generated in the heavy-mass expansion of the determinant provides an additive renormalization of the vacuum-energy/string tension sector, already parametrized by the zero-mode dressing $\exp\left(-\kappa S[C]\right)$; only the sum defines the single physical string tension $\sigma_{phys}=2\sqrt{2}\kappa_{phys}$ in the continuum theory, while $M$ acts as a regulator scale removed in the local limit.
@@ 9.3 New view at the cusp singularity
This mechanism is analogous to the continuum limit in Lattice Gauge Theory, where the bare coupling $g_{0}\to 0$ as the lattice spacing $a\to 0$ to keep physical masses finite. In our induced QCD, the cutoff has a physical and mathematical origin in the regularization of the minimal surface, so the limit $m\to\infty$ (where $\lambda\to 1/(\beta_{0}\log m)$) has a precise mathematical meaning, replacing the formal subtraction of infinities with a finite limiting process.
@@ 10.1 Regularization, renormalizability, and the microscopic definition of Planar QCD
However, the lattice definition comes at a cost: it explicitly breaks the continuous symmetries of four-dimensional Euclidean space, from the rotation group $O(4)$ to subtle topological structures like the Hodge duality of the area tensor. For instance, instanton solutions or Hodge-dual minimal surfaces do not exist on a discrete grid; they only emerge in the continuum limit.
@@
Previous lattice approaches illustrate the sensitivity of this limit. The large-N reduced Eguchi-Kawai model, for example, encountered unphysical fixed points. Similarly, the original Induced QCD proposal [15]—which attempted to induce gauge dynamics via heavy free fermions with vanishing color currents on a lattice—failed because the theory collapsed into the wrong phase rather than asymptotically free QCD. Our Elfin theory is a variation on this theme: heavy free fermions inducing QCD. However, the critical distinction is that our fermions live on a rigid minimal surface rather than a four-dimensional lattice. This geometric constraint keeps the theory non-trivial in the local limit, making it a viable regulator for Planar QCD.
A central question arises regarding the renormalization of the Wilson loop itself. Standard perturbation theory dictates that $W[C]$ requires multiplicative renormalization to remove divergences associated with cusps and self-intersections. However, redefining the loop equation solely to render the coordinate-space Wilson loop finite in the local limit obscures the underlying dynamics.
Physical renormalizability demands only the existence of a universal mass scale $\Lambda_{QCD}$ and the finiteness of observable scattering amplitudes. The Wilson loop $W[C]$ is an intermediate, off-shell quantity. Physical observables involve path integrals over quark trajectories, which are governed by a Brownian measure. A Brownian trajectory is nowhere differentiable; it consists of an infinite density of cusps. Attempting to renormalize the Wilson loop for a smooth contour $C$ is physically artificial because the dominant contributions to the path integral come from fractal paths where the concept of a "cusp angle" is ill-defined.
@@ 10.2 Momentum loop path integral
As we discussed in the previous section, the ordinary Wilson loop is not directly observable, and its singularities are merely intermediate results that will eventually disappear (cancel or integrate out) in observables.
Unlike the standard Feynman-Schwinger worldline representation in coordinate space, which is inherently plagued by the UV cusp singularities of $W[C]$, we utilize the phase-space path integral representation introduced by the author in [26]. In this formulation, the planar quark amplitude factorizes into a universal kinematic Dirac phase-space trace $\mathcal{K}[P]$ and the dynamical gluonic momentum-loop functional $W[P+Q]$.
@@
The phase-space formulation possesses a canonical normalization. The functional measure is normalized by the standard symplectic volume element $\mathcal{D}P\mathcal{D}C^{\prime}/(2\pi)$ per continuous degree of freedom. Consequently, in the absence of the interacting Wilson loop (the free-particle limit, $W[C]=1$), integrating over the loop coordinates $C$ natively yields a functional Dirac delta function, enforcing the conservation of local momentum:
@@
This demonstrates that the phase-space integral is well-defined from standard quantum mechanics, unique modulo the infinite 1D diffeomorphism volume $\text{Vol}(\text{Diff}(S^{1}))$ associated with the simultaneous reparametrization of $P(\tau)$ and $C(\tau)$.Conventionally, this $\text{Diff}(S^{1})$ volume is factored out by imposing a proper-time gauge-fixing via an $\int(C^{\prime})^{2}d\tau$ kinetic term. This specific gauge choice inevitably generates the Feynman-Schwinger representation with its standard Brownian (Wiener) measure in loop space—a measure inherently dominated by everywhere-discontinuous fractal paths that artificially generate the UV cusp singularities. In our framework, we recognize that the Brownian measure is just one possible choice of gauge fixing for the 1D diffeomorphisms. We will deliberately abandon it. Instead, as detailed in Section 11, we fix $\text{Diff}(S^{1})$ by imposing the Virasoro constraint $(f^{\prime})^{2}=0$ on the holomorphic twistors parametrizing the loop velocity. This specific gauge choice entirely avoids the fractal artifacts of the Brownian measure and locks the 1D phase-space kinematics of the quark loop to the 2D internal confining geometry of the minimal surface.
@@ 10.3 The Momentum loop equation
The delta function on the right (which caused all the perturbative singularities in $W[C]$) disappeared in the Fourier integral (see original papers [25, 26, 27] for details).
@@ 10.4 Equivalence of loop diffusion operator to the third Magnus form
where $\Theta(\theta-\tau)$ denotes the periodic theta function on the circle, the trilinear product is transformed into a general triple integral of the momentum derivative $P^{\prime}$. The rigorous mathematical foundation for expanding loop functionals in such iterated integrals is provided by Chen’s theorem [1]. Chen established that the algebra of iterated path integrals separates points in loop space, meaning that the infinite hierarchy of these integrals forms a complete functional basis. This guarantees that our formal expansion in terms of Magnus forms is exhaustive.
Through direct tensor contraction, the kinematic tensor $T^{\alpha\beta\gamma}{}_{\nu}$ implements the Dynkin projector [6] onto the nested double commutator $\hat{D}_{\alpha},[\hat{D}_{\alpha},\hat{D}_{\nu}]]$. Here, $\hat{D}=\partial+A$ represents the covariant derivative operator within the gauge field representation of the Wilson loop. This mapping identifies the associative word structure of the integrated momentum field with its unique Lie algebra representation.
Consequently, by Ree’s Theorem [37] for Free Lie Algebras [38], this operator identically annihilates the symmetric shuffle ideal associated with the triple integral. As the point-split coordinates $\theta_{i}$ collapse to the origin, the periodic step-functions define a cyclically ordered simplex. This unique parametric-invariant Lie projection ensures that the functional converges to the third-order Magnus form $\Omega^{(3)}$. Direct evaluation of this triple integral, mediated by the $T$ projector, confirms this algebraic correspondence and yields the canonical normalization coefficient $1/6$.
@@ 10.5 Algebraic recurrence for the Momentum Loop Equation
The Momentum Loop Equation (MLE) exhibits a significant simplification when the momentum functional is represented as a Magnus-ordered exponential:
@@
where $\hat{X}_{\mu}$ represents the operator position of the string endpoint in spacetime. Expanding this trace in Magnus forms $\Omega^{n}$, we obtain universal tensor coefficients $\mathcal{W}_{\alpha_{1}\dots\alpha_{n}}^{(n)}=\mathrm{tr}\{\hat{X}_{\alpha_{1}}\dots\hat{X}_{\alpha_{n}}\}$. The functional derivative acts algebraically on this structure, bringing down commutator insertions:
@@ 10.6 Exact solution up to $\mathcal{O}(P^{\prime 7})$ and non-analyticity of the Taylor expansion
When substituted into the iterated integrals of the Magnus expansion, the $\delta$-function explicitly inserts the macroscopic gap vector $\Delta P_{\mu}\equiv\Omega^{(1)}_{\mu}$ precisely at the boundaries of the integration domain. Because $\Delta P$ corresponds to the first-order Magnus form, this topological closure modifies the $n$-th order form by subtracting the product $\Omega^{(1)}\Omega^{(n-1)}$. By the integration-by-parts identity of the Magnus expansion (the shuffle relations), this product translates to the sum of all interleaved operator positions. The loop-closing transformation thus yields the modified Magnus form:
@@
Substituting this closed-loop basis into the MLE, we dynamically equate the left-hand side kinematic shuffle products against the right-hand side discrete commutator series. Starting from the perturbative vacuum $\mathcal{W}^{(0)}=1$, the algebraic system yields finite solutions for the lowest-order tensors. The fundamental string tension enters at $\mathcal{W}^{(2)}$, and the kinematic stress generated by the loop derivative constrains the cyclic symmetric pairings of $\mathcal{W}^{(4)}$ and $\mathcal{W}^{(6)}$. (The explicit analytic formulas for these invariant tensors are provided in C for small $n$. Crucially, C also contains the exact finite-dimensional linear algebra diagnostics (Table 2) and the explicit Fredholm inconsistency proof demonstrating the absolute lack of solutions at $\mathcal{W}^{(8)}$. The general algorithmic implementation is provided in [30].)
@@ 10.7 The Reasons for the Failure of the Magnus Expansion for $W[P]$
The general reason for the breakdown of the expansion at this order is that the linear space of $\mathcal{W}^{(8)}$ invariant tensors—which is constrained by the Shuffle Ideal and cyclic symmetry—possesses a significantly smaller matrix rank than the highly specific multilinear stress generated by the crossing Kronecker deltas of $\mathcal{W}^{(4)}\times\mathcal{W}^{(4)}$.
This is not a heuristic observation, but a quantified algebraic contradiction. As detailed in C (and summarized in Table 2), extracting the linear system at the 8th order yields 379 geometric constraints but only 212 available parameters (which exhaustively includes the full $\mathcal{W}^{(8)}$ basis, all residual lower-order free parameters, and all gauge/ideal freedoms). The resulting system explicitly violates the Fredholm alternative condition, yielding a positive rank deficiency ($\Delta\text{rank}=1$) and a non-zero projection of the inhomogeneous stress onto the left-null space of the kinematic matrix.
@@
The physical meaning of this failing Taylor–Magnus expansion is of paramount importance. It demonstrates that the MLE is finite, free of UV divergences, and expandable in a functional Taylor series up to $\mathcal{O}(P^{\prime 6})$. However, the geometric contradiction at the 8th order proves that the solution $W[P]$ is not an analytic functional of the bounding loop $P(\theta)$. It cannot be indefinitely expanded in a continuous Taylor–Magnus series.
@@
governing a scalar functional $W[P]$. The number of independent tensor structures available for a scalar functional grows slower than those required for a free-index vector equation (growing approximately four times slower in $d=4$). This is why, at some critical level of the Magnus expansion, there will inevitably be more vector equations than available scalar parameters.
This is a foundational feature of the loop equation: it descends directly from the vector Yang–Mills equations of motion, which dictate that a local vector loop equation (evaluated at a fixed point on the contour) must be satisfied. Resolving this overdetermined system demands a mathematical “conspiracy” among the internal degrees of freedom within $W[P]$. The generic Taylor–Magnus expansion simply lacks the internal geometric structure to orchestrate this conspiracy.
(Interestingly, the same vector-versus-scalar mismatch is fatal to all naive bosonic string theories of QCD: the loop equations are vector equations, but the Nambu–Goto string Laplace operator is a scalar, meaning a scalar area-derivative constraint fundamentally cannot capture the full vector loop equation without an anomaly.)
However, our theory natively provides this required conspiracy! In the previous sections, we derived vector equations for the Elfin determinant on the Hodge-dual minimal surface, and these vector equations are satisfied across all vector components. This property is made possible by Hodge duality and the planar factorization of the Majorana determinant on this Hodge dual surface.
The physical reasons for this non-analyticity will become fully apparent in later sections, when we construct the continuum solution and reduce it to the twistor string path integral. In that formulation, the formal Taylor expansion in $P(\theta)$ corresponds to the perturbative expansion of the exponential phase factor:
@@
For such a functional Taylor series to converge, the boundary twistors $\lambda(z)$ and $\mu(z)$ would need to be bounded. The algebraic contradiction at the 8th order proves this is not the case; the path integral is not convergent in a perturbative Taylor expansion.
As we will discuss in detail when deriving the twistor string path integral, the integration over the momentum loops is not dominated by smooth, small $P$ variations at all. When integrated alongside the Dirac target-space product $\mathrm{tr}\,\hat{\mathcal{T}}\exp\left(-\mathrm{i}\int d\theta\,\gamma_{\mu}P_{\mu}(\theta)+m_{0}\right)$, we encounter ultra-local integrals. At every angle $\theta$, we integrate out the local value $P(\theta)$. Because there are no worldline kinetic derivatives $\partial_{\theta}P$ in the action, the path integral completely factorizes point-by-point. As a result, the dominant trajectories are neither continuous nor dominated by small $P$ variations. We therefore conclude that the MLE is completely finite and anomaly-free, but its true continuum solution is inherently non-perturbative and cannot be Taylor-expanded beyond the 6th order.